OpenStudy (anonymous):

Convert -270 to radian measure.

OpenStudy (psymon):

Degrees to radians - \[degrees*\frac{ \pi }{ 180 }\]

OpenStudy (anonymous):

-270 degrees, right?

OpenStudy (anonymous):

\[180^{o} = \pi^c\] this is the conversion rule , now you can use unitary method to find your answer. if 180 degrees = 22/7 radians 1 degree = (22/7) / 180 270 degrees = 270* {(22/7) / 180]

OpenStudy (anonymous):

270 degrees is three fourths of a circle. In radians, the circle is 2pi. Three fourths of that is 3pi/2. The opposite of that (measuring clockwise) would be -3pi/2.

OpenStudy (anonymous):

Radian measures are almost always given in multiples of pi.

OpenStudy (anonymous):

So the final answer would be -3pi/2?

OpenStudy (skullpatrol):

yes

OpenStudy (skullpatrol):

$$\frac{-3\pi}{2}$$

OpenStudy (anonymous):

Thank you!

OpenStudy (skullpatrol):

thanks for learning